OpenStudy (anonymous):

sin(5pi/12)+sin(pi/12) what is the exact value using sum to product identities? Please help!

OpenStudy (***[isuru]***):

hi, \[\sin C + \sin D = 2\sin( \frac{ C +D }{ 2 })\cos(\frac{ C-D }{ 2 })\] so... we can write ur answer... \[\sin \frac{ 5 \pi }{ 12} + \sin \frac{ \pi }{12 } = 2\sin(\frac{ 5 \pi + \pi }{ 12 \times 2 })\cos( \frac{ 5 \pi - \pi }{12 \times 2 })\]\[2 \sin \frac{ 6 \pi }{24 }\cos \frac{ 4 \pi }{ 24 }\]\[2 \sin \frac{ \pi }{ 4 }\cos \frac{ \pi }{6 }\] now... \[\sin \frac{ \pi }{ 4 } = \frac{ 1 }{ \sqrt{2} } \ and \ \cos \frac{ \pi }{ 6 } =\frac{ \sqrt{3} }{ 2 }\] which means...\[\sin \frac{ 5 \pi }{12 } + \cos \frac{ \pi }{12 } = 2 \times \frac{ 1 }{\sqrt{2} } \times \frac{ \sqrt{3} }{ 2 }\] \[= \frac{ \sqrt{3} }{ \sqrt{2}} = \frac{ \sqrt{6} }{ 2}\] Hope this will help ya!!

OpenStudy (anonymous):

yes thank you ! i didnt add the 12+12 at the bottom so keep getting wrong/

OpenStudy (***[isuru]***):

hah.. but now u got it correct!! that's the important part!!